# Signed Higher-Order Interactions for Brain Disorder Diagnosis via Multi-Channel Transformers

### v1

@@ Abstract
Accurately characterizing the higher-order interactions of brain regions and effectively extracting the interpretable higher-order organizational patterns in Functional Magnetic Resonance Imaging (fMRI) data are crucial for the diagnosis of brain diseases. However, current graph models based on deep learning mainly focus on pairwise patterns, as well as triadic patterns within brain while overlooking more higher-order patterns with signs, limiting an integrated understanding of brain-wide communication. To address these challenges, we propose HOI-Brain (Higher-Order Interaction in Brain Network), a novel computational framework that enables the utilization of signed higher-order interactions and signed organizational patterns in fMRI data for the diagnosis of brain diseases. Specifically, we present a new calculation of co-fluctuations based on Multiplication of Temporal Derivatives to detect higher-order interactions with adequate temporal resolution. Next, we further distinguish positively and negatively synergistic higher-order interactions and encode them in signed weighted simplicial complexes, which can offer detailed insights into the communication within the brain. Moreover, founded upon Persistent Homology theory, two distinct filtration processes are employed in weighted simplicial complexes of the brain to extract signed higher-dimensional neural organisations from a spatiotemporal perspective. Finally, a multi-channel brain Transformer architecture is proposed to holistically integrate information from heterogeneous topological features. Comprehensive experiments across Alzheimer’s disease, Parkinson’s syndrome, and autism spectrum disorder datasets demonstrate the superiority, effectiveness, and interpretability of our framework. The extracted key brain regions and higher-order organizational patterns have been demonstrated to align reliably with domain knowledge from the neuroscience scientific literature, thereby providing direct and meaningful insights. Our code is avaiable.
@@
journal: Medical Image Analysis
@@ 1 Introduction
Functional Magnetic Resonance Imaging (fMRI) is one of the key tools in neuroscience to discover potential neuroimaging biomarkers for the automated diagnosis of brain disorders, including Alzheimer’s disease (AD), autism spectrum disorder (ASD), and Parkinson’s disease (PD) (Lindquist, 2008; Fornito et al., 2016). Specifically, resting state fMRI measures functional connectivity between brain regions via blood-oxygen-level-dependent (BOLD) signals. This process naturally models the brain as a network, defining regions of interest (ROIs) as nodes and functional connections between ROIs as edges (Sporns, 2010; Bullmore and Sporns, 2009). Abnormalities in this network at the whole brain level are expected to yield disease-related biomarker signatures (Wang et al., 2021).
@@
A significant number of graph data mining methodologies are then applied to such brain networks for brain disorder diagnosis, with the goal of understanding and analysing the elements and interactions of neurological systems from a network perspective. Examples include graph neural network (GNN)-based models (Xia et al., 2025), brain Transformer-based models (Kan et al., 2022b), hypergraph neural network (HGNN)-based models (Wang et al., 2024a), and persistent homology (PH)-based models (Bian et al., 2023). Although these studies have demonstrated the considerable potential of network based approaches in elucidating the complexities of the brain and diagnosing brain disorders, graph analysis of brain networks remains an emerging field in its infancy (Fotiadis et al., 2024).
@@
Recently, graph neural network (GNN) techniques have been employed to analyse brain networks characterised by pairwise interactions. These methods facilitate the extraction of potential topological features that are instrumental in diagnosing brain disorders (Luo et al., 2024). Several GNN-based models specifically designed for brain networks—including GroupINN (Yan et al., 2019), BrainGNN (Li et al., 2021), FBNetGen (Kan et al., 2022a), BPI-GNN (Zheng et al., 2024), and ASD-HNet (Luo et al., 2025)—have been proposed and have achieved promising diagnostic performance. It is generally accepted that GNN-based models use a message passing mechanism in which the embedding of a brain ROI is updated by aggregating information from its neighbouring ROIs, thereby learning a discriminative graph-level representation of the brain connectivity network. However, these methods are limited by the underlying assumption that interactions between nodes are strictly pairwise. Growing evidence indicates that the complexity of brain-region interactions extends far beyond pairwise connections, with widespread co-fluctuations occurring in groups of nodes that evolve over time (Battiston et al., 2020; Chelaru et al., 2021). Consequently, the limited expressiveness of traditional GNNs prevents them from capturing higher-order interactions in brain networks, and the topological features they extract are typically regarded as lower-order.
@@
Transformer-based brain network models have achieved tremendous success in diagnosing brain disorders, thanks to their ability to capture global patterns. Representative examples include Graph Transformer (Ying et al., 2021), Brain Network Transformer (Kan et al., 2022b), TSEN (Hu et al., 2023), and Long-range Brain Transformer (Yu et al., 2024). These models naturally construct fully connected graphs and, through a powerful global attention mechanism, adaptively learn pairwise interaction relationships for brain disorder diagnosis. However, they typically use either the raw time series features of brain regions or the functional connectivity matrix as node-level input, thereby overlooking significant higher-order topological information among ROIs.
@@
Several studies have focused on representing higher-order interactions (HOIs) in brain networks with more sophisticated models—primarily hypergraphs, thereby improving diagnostic performance for brain diseases (Jie et al., 2016; Xiao et al., 2019; Wang et al., 2022; Hao et al., 2023; Wang et al., 2024a). Although HGNN-based models have shown promising results by characterising HOIs via hyperedges generated through k-nearest-neighbour or k-hop neighbourhoods, this node-centric construction scheme conflicts with the notion of group dependence, as it fails to capture genuine simultaneous interactions among groups of ROIs. Moreover, it has been demonstrated that higher-order features extracted by HGNN-based models can obscure their relationship to specific topological or neurobiological phenomena (Kim et al., 2024; Su et al., 2024).
@@
To extract more interpretable higher-order topological features for brain-disorder diagnosis, recent studies have adopted persistent homology (PH), a topological approach capable of reconstructing HOI structures and delivering state-of-the-art performance in characterising brain topological profiles (Hyekyoung et al., 2011; Sizemore et al., 2018). However, most investigations of higher-order structures in brain networks have focused on 0-dimensional (connected components) and 1-dimensional (cycles) topological profiles formed by nodes and edges; these dimensions do not link HOIs in the brain to higher-order organisations that would provide a truly higher-dimensional perspective (Talesh Jafadideh and Mohammadzadeh Asl, 2022; Bian et al., 2024; Bhattacharya et al., 2025). To address this limitation, researchers have recently leveraged synchronisation phenomena to construct more accurate simplicial complexes for modelling HOIs and have used PH to capture 1-dimensional cycles that reflect triplet interactions (Santoro et al., 2024). Comprehensive analyses demonstrate that methods incorporating inferred higher-order interactions among three ROIs outperform traditional pairwise approaches, thereby offering new insights into the higher-order organisation of fMRI time-series data. Nevertheless, evidence shows that triplet interactions can be decomposed into linear combinations of pairwise interactions if they are linearly decomposable, implying that some triplets may not represent genuine higher-order phenomena (Delabays et al., 2025; Neuhäuser et al., 2020). Consequently, considering even higher-order interactions and organisational structures could yield deeper understanding of brain function. Moreover, existing measures of instantaneous co-fluctuation rely on extended Pearson correlation, which lacks the temporal resolution required to detect spatiotemporal interactions among groups of regions (Shine et al., 2015). Finally, the role of signed HOIs—potentially informative for neurological conditions—has been largely overlooked in these studies.
@@
To address these methodological challenges, we propose an innovative computational framework, HOI-Brain (Higher-Order Interactions in Brain), that enables a comprehensive analysis of HOIs in fMRI data (Fig. 1). First, we introduce a novel metric—Multiplication of Temporal Derivatives (MTD)—to quantify dynamic functional co-fluctuations among groups of ROIs. MTD computes element-wise products of the temporal derivatives of blood-oxygen-level-dependent (BOLD) signals, yielding instantaneous co-fluctuation magnitudes for k-node interactions. This approach more reliably identifies genuine higher-order neural interactions than Pearson’s correlation. Second, adopting a spatiotemporal analytic perspective, we model brain networks as weighted simplicial complexes that link instantaneous co-fluctuation of k node interactions in the brain to the weight of k-simplex, which can accommodate more information about interactions among multi-ROIs compared to pairwise networks. We further extract positively and negatively synergistic interactions from the weighted simplicial complexes to construct two signed weighted simplicial complexes, respectively. Positively synergistic interactions reflect multiple brain regions that exhibit simultaneous activation at a given moment relative to the preceding one, while negatively synergistic interactions indicate that these regions collectively exhibit inhibition at the current moment compared to the prior moment, which can offer detailed information about complex coordination and communication within the brain, thus improving the effectiveness of brain disease diagnosis. We further explore the influence of signed higher-order interactions (HOIs) using the self-attention mechanism. Through two distinct filtration processes based on Persistent Homology theory, we systematically extract four types of higher-order topological features: (1) positive and negative quadruplet-level interaction signatures that capture irreducible higher-order neural coordination patterns and (2) positive and negative two-dimensional void descriptors that characterize the intrinsic geometric organization of neural activity from a higher-dimensional manifold perspective. Third, given the significance of both lower-order interactions and higher-order interactions in brain network analysis tasks, we develop a multi-channel brain network Transformer to synergistically integrate lower-order edge features with the above four types of higher-order topological invariants. Specifically, we inject higher-order information into a multi-channel Transformer and use an orthonormal clustering readout operation based on self-supervised soft clustering and orthonormal projection to learn distinguishable cluster-aware lower- and higher-order node embeddings. Then we use an attention mechanism to adaptively fuse the features encoded by different channels. Finally, we utilize MLP as prediction head to classify the different levels of topological features extracted from the multi-channel brain network Transformer.
@@
Figure 1: Overall framework of HOI-Brain. Individual fMRI data are transformed into $N$ original fMRI signals through a preprocessing pipeline. A novel metric - Multiplication of Temporal Derivatives (MTD) - is quantified dynamic functional co-fluctuations of group ROIs. Then, at each timepoint t, some instantaneous k-order co-fluctuation time series are encoded into two signed weighted simplicial complexes respectively. Persistent Homology are applied to analyze the two signed weighted simplicial complexes at each t. Five feature matrices are generated by extracting all edges, signed good quadruplets, and signed 2D voids. These matrices are temporally averaged across all timepoints to stabilize feature feature representations. By incorporating the lower-order and higher-order feature, a multi-channel brain network transformer is using to produce the final classification.
@@
Experiments on Alzheimer’s disease, autism spectrum disorder, and Parkinson’s disease datasets have demonstrated its superiority over 20 state-of-the-art baselines. The present study employs an ablation approach to demonstrate the significance of higher-order topological features, the efficacy of distinct components of the proposed model, and the superiority of quadruplet interactions over triplet interactions in enhancing the diagnostic efficacy of brain diseases. Next, by visualizing the attention scores in the self-attention mechanism, the importance of signed information and channel information was identified. This finding indicates that the integration of positive and negative information, in conjunction with higher-order and low-order interactions within brain networks, can remarkably enhance diagnostic performance in a range of brain disease diagnostic tasks. Furthermore, by visualising the self-attention maps in the multi-channel brain network Transformer, interaction patterns among disease-related key brain regions and key interactions between ROIs were identified from a more comprehensive brain network perspective. In addition, we uncover some higher-order organizational patterns associated with specific stages of disease progression based on our methods.
@@ 2.1 Graph models based on deep learning
A significant number of graph models are then applied to brain networks for the purpose of brain disorder diagnosis, with the objective being to understand and analyse the elements and interactions of neurological systems from a network perspective. Graph neural networks (GNNs) have been demonstrated to be particularly effective in modelling brain connectomes, due to their ability to effectively capture subtle, latent representations and nonlinear relationships in graph-structured data. Several GNN-based models, specifically designed for brain networks, have been proposed and have achieved promising diagnostic performance. For instance, GroupINN (Yan et al., 2019) incorporates the concept of node grouping into the neural network and designs a random-walk-based variant of graph convolutional layer. BrainGNN (Li et al., 2021) designs novel ROI-aware graph convolutional (Ra-GConv) layers that leverage the topological and functional information of fMRI. FBNETGEN (Kan et al., 2022a) utilizes a task-aware GNN-based framework for fMRI analysis via functional brain network generation, which generates the brain connectivity matrices and predicts clinical outcomes simultaneously from fMRI BOLD signal series. BPI-GNN (Zheng et al., 2024) analyzes functional magnetic resonance images (fMRI) by leveraging the well-known prototype learning method. ASD-HNNet (Luo et al., 2025) designs a hybrid neural network model for ASD identification, which hierarchically extracts features from functional brain networks at threescales: local ROI scale, community scale, and global representation scale. However, these methods are essentially based on a two-way network to model pairwise interactions and mine low-order topological features for the diagnosis of brain disorders, overlooking the influence of HOIs in the brain.
@@
Additionally, with advancements in large language models, transformer-based methods are increasingly being integrated into brain graph structures. Graph transformer (Ying et al., 2021) can be applied to brain networks to learn pairwise connection strengths among ROIs across individuals. Then, Brain Network Transformer (Kan et al., 2022b) leverages the unique properties of brain network data to fully unleash the power of Transformer-based models for brain network analysis, which avoids the costly computations of eigenvalues or eigenvectors. TSEN (Hu et al., 2023) introduced snowball graph convolution as position embedding in Transformer structure, which was a simple yet effective method for capturing local patterns of brain naturally. Long-range Brain Transformer (Yu et al., 2024) injects the long-range embeddings into a transformer framework and integrate both short-range and long-range dependencies between ROIs using the self-attention mechanism. These works ignore significant high-order topological information among ROIs.
@@
Some studies have been made in constructing HOIs in the brain to improve the performance of brain disorder diagnosis based on hypergraphs. For example, a novel framework (Jie et al., 2016) is proposed to estimate the hyperconnectivity network of brain functions for the diagnosis of brain disease. MHL-Hypergraph (Xiao et al., 2019) proposes a multi-hypergraph learning based method to compute a unified hypergraph similarity matrix from multi-paradigm fMRI data to represent an FCN for each subject. An evolving hypergraph convolutional network (Wang et al., 2022) for the dynamic hyperbrain network is proposed, which adds the attention mechanism to further enhance the ability of representation learning. To consider the temporal characteristics of longitudinal data, a weighted hypergraph convolution network (WHGCN) (Hao et al., 2023) is designed to use the internal correlations among different time points and leverage high-order relationships between subjects for AD detection. CcSi-MHAHGEL (Wang et al., 2024a) introduced a new hypergraph convolution network framework to extract multiatlas-based FCN embeddings with application to multisite ASD identification, where hyperedge-aware HGCN was developed to capture complex high order information in brain network. These studies primarily use k-nearest neighbour or k-hop neighbour constructions to generate hypergraphs, which are inconsistent with the definition of group dependence in brain regions. Furthermore, higher-order features extracted using hypergraph neural networks (HGNNs) are frequently regarded as being less interpretable.

### v2

@@ Abstract
Accurately characterizing the higher-order interactions of brain regions and effectively extracting the interpretable higher-order organizational patterns from Functional Magnetic Resonance Imaging (fMRI) data are crucial for the diagnosis of brain diseases. However, current graph models based on deep learning mainly focus on pairwise patterns, as well as triadic patterns within brain while overlooking more higher-order patterns with signs, limiting an integrated understanding of brain-wide communication. To address these challenges, we propose HOI-Brain (Higher-Order Interaction in Brain Network), a novel computational framework that enables the utilization of signed higher-order interactions and signed organizational patterns in fMRI data for the diagnosis of brain diseases. Specifically, we present a new calculation of co-fluctuations based on Multiplication of Temporal Derivatives to detect higher-order interactions with adequate temporal resolution. Next, we further distinguish positively and negatively synergistic higher-order interactions and encode them in signed weighted simplicial complexes, which can offer detailed insights into the communication within the brain. Moreover, founded upon Persistent Homology theory, two distinct filtration processes are employed in weighted simplicial complexes of the brain to extract signed higher-dimensional neural organisations from a spatiotemporal perspective. Finally, a multi-channel brain Transformer architecture is proposed to holistically integrate information from heterogeneous topological features. Comprehensive experiments across Alzheimer’s disease, Parkinson’s syndrome, and autism spectrum disorder datasets demonstrate the superiority, effectiveness, and interpretability of our framework. The extracted key brain regions and higher-order organizational patterns have been demonstrated to align reliably with domain knowledge from the neuroscience scientific literature, thereby providing direct and meaningful insights. Our code is avaiable.
@@
journal: Elsevier
@@ 1 Introduction
Functional magnetic resonance imaging (fMRI) represents a pivotal instrument in the field of neuroscience, serving to identify potential neuroimaging biomarkers. These biomarkers are instrumental in the automated diagnosis of various brain disorders, including Alzheimer’s disease (AD), autism spectrum disorder (ASD), and Parkinson’s disease (PD) (Lindquist, 2008; Fornito et al., 2016). Specifically, resting state fMRI measures functional connectivity between brain regions via blood-oxygen-level-dependent (BOLD) signals. This process naturally models the brain as a network, defining regions of interest (ROIs) as nodes and functional connections between ROIs as edges (Sporns, 2010; Bullmore and Sporns, 2009). Abnormalities in this network at the whole-brain level are hypothesised to yield disease-related biomarker signatures (Wang et al., 2021).
@@
A significant number of graph data mining methodologies are then applied to such brain networks for brain disorder diagnosis, with the goal of understanding and analysing the elements and interactions of neurological systems from a network perspective. Examples of such models include graph neural network (GNN)-based models (Xia et al., 2025), brain Transformer-based models (Kan et al., 2022b), hypergraph neural network (HGNN)-based models (Wang et al., 2024a), and persistent homology (PH)-based models (Bian et al., 2023). Despite the evidence from these studies demonstrating the considerable potential of network-based approaches to elucidating the complexities of the brain and diagnosing brain disorders, graph analysis of brain networks remains an emerging field in its infancy (Fotiadis et al., 2024).
@@
In recent developments, graph neural network (GNN) techniques have been employed to analyse brain networks characterised by pairwise interactions. The utilisation of these methodologies enables the extraction of potential topological features, which are instrumental in the diagnosis of brain disorders (Luo et al., 2024). A number of GNN-based models have been proposed for brain networks, including GroupINN (Yan et al., 2019), BrainGNN (Li et al., 2021), FBNetGen (Kan et al., 2022a), BPI-GNN (Zheng et al., 2024), and ASD-HNet (Luo et al., 2025). These models have demonstrated the potential to achieve favourable diagnostic performance. It is widely accepted that GNN-based models utilise a message passing mechanism, whereby the embedding of a brain ROI is updated by aggregating information from its neighbouring ROIs. This process facilitates the learning of a discriminative graph-level representation of the brain connectivity network. However, these methods are limited by the underlying assumption that interactions between nodes are strictly two-way. Recent studies have indicated a growing body of evidence that suggests the complexity of brain-region interactions extends far beyond the established pairwise connections. There is now a growing consensus that widespread co-fluctuations occur in groups of nodes that evolve over time (Battiston et al., 2020; Chelaru et al., 2021). Consequently, the limited expressiveness of traditional GNNs prevents them from capturing higher-order interactions in brain networks, and the topological features they extract are typically regarded as lower-order.
@@
Transformer-based brain network models have achieved considerable success in the diagnosis of brain disorders, due to their capacity to capture global patterns. Representative examples include Graph Transformer (Ying et al., 2021), Brain Network Transformer (Kan et al., 2022b), TSEN (Hu et al., 2023), and Long-range Brain Transformer (Yu et al., 2024). It is evident that these models naturally construct fully connected graphs and, through a powerful global attention mechanism, adaptively learn pairwise interaction relationships for brain disorder diagnosis. However, the prevailing approach involves the utilisation of either the raw time series features of brain regions or the functional connectivity matrix as node-level input, thereby overlooking significant higher-order topological information among ROIs.
@@
A number of studies have concentrated on the representation of higher-order interactions (HOIs) in brain networks using more sophisticated models, with a particular focus on hypergraphs. This approach has been shown to enhance diagnostic performance for brain diseases, as evidenced by citations in the literature (Jie et al., 2016; Xiao et al., 2019; Wang et al., 2022; Hao et al., 2023; Wang et al., 2024a). Despite the encouraging outcomes demonstrated by HGNN-based models in characterising HOIs through hyperedges generated via k-nearest neighbour or k-hop neighbourhoods, this node-centric construction scheme appears to be incompatible with the concept of group dependence. This is due to its inability to capture authentic simultaneous interactions among groups of ROIs. Furthermore, it has been demonstrated that higher-order features extracted by HGNN-based models can obscure their relationship to specific topological or neurobiological phenomena (Kim et al., 2024; Su et al., 2024).
@@
In order to extract higher-order topological features that are more readily interpretable for the purpose of brain-disorder diagnosis, recent studies have adopted persistent homology (PH), a topological approach capable of reconstructing HOI structures and delivering state-of-the-art performance in characterising brain topological profiles (Hyekyoung et al., 2011; Sizemore et al., 2018). However, the majority of investigations into higher-order structures in brain networks have focused on 0-dimensional (connected components) and 1-dimensional (cycles) topological profiles formed by nodes and edges. These dimensions do not link HOIs in the brain to higher-order organisations that would provide a truly higher-dimensional perspective (Talesh Jafadideh and Mohammadzadeh Asl, 2022; Bian et al., 2024; Bhattacharya et al., 2025). In order to address the aforementioned limitation, researchers have recently employed synchronisation phenomena to construct more accurate simplicial complexes for modelling HOIs. Furthermore, they have utilised PH to capture 1-dimensional cycles that reflect triplet interactions (Santoro et al., 2024). Comprehensive analyses demonstrate that methods incorporating inferred higher-order interactions among three ROIs outperform traditional pairwise approaches, thereby offering new insights into the higher-order organisation of fMRI time-series data. Nonetheless, the extant evidence indicates that triplet interactions can be decomposed into linear combinations of pairwise interactions, provided that said interactions are linearly decomposable. This finding suggests that some triplets may not in fact represent genuine higher-order phenomena (Delabays et al., 2025; Neuhäuser et al., 2020). Consequently, consideration of higher-order interactions and organisational structures may offer a more profound comprehension of brain function. Furthermore, extant measures of instantaneous co-fluctuation are contingent on extended Pearson correlation, a method that is deficient in temporal resolution when it comes to detecting spatiotemporal interactions among groups of regions (Shine et al., 2015). Finally, the role of signed HOIs, which have the potential to provide valuable diagnostic information in cases of neurological conditions, has been largely overlooked in these studies.
@@
In order to address the methodological challenges identified, a novel computational framework has been proposed. This framework, designated HOI-Brain (Higher-Order Interactions in Brain), facilitates a comprehensive analysis of HOIs in fMRI data 1. Firstly, a novel metric – Multiplication of Temporal Derivatives (MTD) – is introduced with the aim of quantifying dynamic functional co-fluctuations among groups of ROIs. MTD performs element-wise product calculations on the temporal derivatives of blood-oxygen-level-dependent (BOLD) signals, thereby yielding instantaneous co-fluctuation magnitudes for k-node interactions. This approach has been demonstrated to offer a more reliable means of identifying genuine higher-order neural interactions in comparison to Pearson’s correlation. Secondly, adopting a spatiotemporal analytic perspective, we propose a model for brain networks as weighted simplicial complexes. These complexes link instantaneous co-fluctuation of k-node interactions in the brain to the weight of the k-simplex. This facilitates the accommodation of more information about interactions among multiple regions of interest (ROIs) in comparison to pairwise networks. The following step involves the extraction of both positive and negative synergistic interactions from the weighted simplicial complexes in order to construct two signed weighted simplicial complexes, respectively. Positively synergistic interactions are indicative of multiple brain regions that exhibit simultaneous activation at a given moment relative to the preceding one, while negatively synergistic interactions indicate that these regions collectively exhibit inhibition at the current moment compared to the prior moment. Collectively, these interactions offer detailed information about complex coordination and communication within the brain, thus improving the effectiveness of brain disease diagnosis. The present study explores the influence of signed higher-order interactions (HOIs) using the self-attention mechanism. The extraction of four types of higher-order topological features is achieved through the implementation of two distinct filtration processes, which are based on Persistent Homology theory. The first type of interaction signature is positive and negative quadruplet-level interaction signatures. These capture irreducible higher-order neural coordination patterns. The second type of interaction signature is positive and negative two-dimensional void descriptors. These characterise the intrinsic geometric organisation of neural activity from a higher-dimensional manifold perspective. Thirdly, given the significance of both lower-order interactions and higher-order interactions in brain network analysis tasks, a multi-channel brain network Transformer is developed to synergistically integrate lower-order edge features with the above four types of higher-order topological invariants. Specifically, higher-order information is injected into a multi-channel Transformer, and an orthonormal clustering readout operation is employed. This operation is based on self-supervised soft clustering and orthonormal projection, and it is used to learn distinguishable cluster-aware lower- and higher-order node embeddings. Subsequently, an attention mechanism is employed to facilitate the adaptive fusion of features encoded by disparate channels. In conclusion, the Multi-Layer Perceptron (MLP) is employed as the prediction head to categorise the various levels of topological features that have been extracted from the multi-channel brain network Transformer.
@@
Figure 1: Overall framework of HOI-Brain. Individual fMRI data are transformed into $N$ original fMRI signals through a preprocessing pipeline. A novel metric - Multiplication of Temporal Derivatives (MTD) - is used to quantify dynamic functional co-fluctuations of group ROIs. Then, at each timepoint t, some instantaneous k-order co-fluctuation time series are encoded into two signed weighted simplicial complexes respectively. Persistent Homology is applied to analyze the two signed weighted simplicial complexes at each t. Five feature matrices are generated by extracting all edges, signed good quadruplets, and signed 2D voids. These matrices are temporally averaged across all timepoints to stabilize feature representations. By incorporating the lower-order and higher-order feature, a multi-channel brain network Transformer is used to produce the final classification.
@@
Experiments on Alzheimer’s disease, autism spectrum disorder, and Parkinson’s disease datasets have demonstrated its superiority over 20 state-of-the-art baselines. The present study employs an ablation approach to demonstrate the significance of higher-order topological features, the efficacy of distinct components of the proposed model, and the superiority of quadruplet interactions over triplet interactions in enhancing the diagnostic efficacy of brain diseases. In the subsequent stage of the research, the importance of signed information and channel information was identified by means of visualising the attention scores in the self-attention mechanism. This finding indicates that the integration of positive and negative information, in conjunction with higher-order and low-order interactions within brain networks, can remarkably enhance diagnostic performance in a range of brain disease diagnostic tasks. By visualising the self-attention maps in the multi-channel brain network Transformer, interaction patterns among disease-related key brain regions and key interactions between ROIs were identified from a more comprehensive brain network perspective. Furthermore, we have identified higher-order organisational patterns associated with specific stages of disease progression, based on the methods employed.
@@ 2.1 Graph models based on deep learning
A significant number of graph models are then applied to brain networks for the purpose of brain disorder diagnosis, with the objective being to understand and analyse the elements and interactions of neurological systems from a network perspective. Graph neural networks (GNNs) have been demonstrated to be particularly effective in modelling brain connectomes, due to their ability to effectively capture subtle, latent representations and nonlinear relationships in graph-structured data. A number of GNN-based models, specifically designed for brain networks, have been proposed, and these have achieved a promising diagnostic performance. For instance, GroupINN (Yan et al., 2019) incorporates the concept of node grouping into the neural network and designs a random-walk-based variant of graph convolutional layer. BrainGNN (Li et al., 2021) has developed a novel ROI-aware graph convolutional (Ra-GConv) layer that utilises the topological and functional information of fMRI. FBNETGEN (Kan et al., 2022a) employs a task-aware GNN-based framework for fMRI analysis via functional brain network generation, which generates the brain connectivity matrices and predicts clinical outcomes simultaneously from fMRI BOLD signal series. BPI-GNN (Zheng et al., 2024) utilises the prototype learning method to analyse functional magnetic resonance images (fMRI). ASD-HNNet (Luo et al., 2025) is a hybrid neural network model for the identification of autism spectrum disorder (ASD). This model extracts features from functional brain networks at three distinct levels: local ROI, community, and global representation. However, these methodologies are fundamentally predicated on a two-way network to model pairwise interactions and mine low-order topological features for the diagnosis of brain disorders, overlooking the influence of HOIs in the brain.
@@
Furthermore, with the advent of sophisticated large language models, Transformer-based methodologies are being increasingly incorporated into brain graph structures. The Graph Transformer (Ying et al., 2021), a sophisticated machine learning algorithm, can be applied to brain networks to learn the strength of the connections between regions of interest (ROIs) across individuals. Subsequently, the Brain Network Transformer (BNT) (Kan et al., 2022b) employs the distinctive characteristics inherent in brain network data to maximise the efficacy of Transformer-based models for brain network analysis, thereby circumventing the necessity for time-consuming computations of eigenvalues or eigenvectors. TSEN (Hu et al., 2023) pioneered the incorporation of snowball graph convolution as position embedding within the Transformer structure, a methodology that has been proven to be both straightforward and efficacious in the context of capturing local patterns of brain activity in a natural manner. Long-range Brain Transformer (Yu et al., 2024) injects the long-range embeddings into a Transformer framework, integrating both short-range and long-range dependencies between regions of interest (ROIs) using the self-attention mechanism. It is evident that these works disregard substantial higher-order topological information among ROIs.
@@
A number of studies have been conducted on the construction of HOIs in the brain with a view to enhancing the performance of brain disorder diagnosis based on hypergraphs. For instance, a novel framework is proposed to estimate the hyper-connectivity network of brain functions for the diagnosis of brain disease (Jie et al., 2016). MHL-Hypergraph (Xiao et al., 2019) proposes a multi-hypergraph learning-based method to compute a unified hypergraph similarity matrix from multi-paradigm fMRI data to represent an FCN for each subject. An evolving hypergraph convolutional network (Wang et al., 2022) for the dynamic hyper-brain network is proposed, which adds the attention mechanism to further enhance the ability of representation learning. In order to take the temporal characteristics of longitudinal data into consideration, a weighted hypergraph convolution network (WHGCN) has been designed to utilise the internal correlations among different time points and to leverage higher-order relationships between subjects for the purpose of AD detection (Hao et al., 2023). CcSi-MHAHGEL (Wang et al., 2024a) pioneered a novel hypergraph convolution network framework for extracting multiatlas-based FCN embeddings, with the objective of facilitating multisite ASD identification. In this framework, hyperedge-aware HGCN was developed to capture complex high-order information in brain networks. The majority of these studies employ k-nearest neighbour or k-hop neighbour constructions in order to generate hypergraphs. However, it has been demonstrated that these constructions are inconsistent with the definition of group dependence in brain regions. Moreover, higher-order features extracted using hypergraph neural networks (HGNNs) are frequently regarded as being less interpretable.